OpenStudy (anonymous):

The volume of a fish tank is 40 cubic feet. If the density is 0.4fish over feet cubed, how many fish are in the tank? 16 4 100 10

OpenStudy (anonymous):

@Directrix help plz

Directrix (directrix):

I'm not sure about density. My instinct is to set up a proportion. .4 fish is to one cubic foot as total fish in the tank is to 40 cubic feet. .4/1 = F/40 where F is the total number of fish in the tank @luka1998 Cross multiply and solve for F. Let's see if the answer makes sense.

Directrix (directrix):

.4/1 = F/40 F*1 = .4 * 40 F = .4 * 40 F = ? @luka1998

OpenStudy (anonymous):

160

Directrix (directrix):

@luka1998 I think you missed the decimal. Try again. 160 is too many fish.

OpenStudy (anonymous):

16

Directrix (directrix):

Okay. That is what I got. Let's see if it makes sense. 16 fish in a 40 gallon tank would be a density of 16/40 = .4 fish per cubic foot. So, that sounds correct.

OpenStudy (anonymous):

@Directrix thx

Directrix (directrix):

You are welcome.

OpenStudy (shaw74):

answer is 16 to be clear!!!!